# A Compressive-Expressive Communication Framework for Compositional Representations

### v3

@@ Abstract
Compositional generalization—the ability to interpret novel combinations of familiar elements—is a hallmark of human cognition and language. Despite recent advances, deep neural networks still struggle to acquire this property reliably. In this work, we introduce CELEBI (Compressive-Expressive Language Emergence through a discrete Bottleneck and Iterated learning), a novel self-supervised framework for inducing compositionality in learned representations from pre-trained models, through a reconstruction-based communication game between a sender and a receiver. Building on theories of language emergence, we integrate three mechanisms that jointly promote compressibility, expressivity, and efficiency in the emergent language. First, interactive decoding incentivizes intermediate reasoning by requiring the receiver to produce partial reconstructions after each symbol. Second, a reconstruction-based imitation phase, inspired by iterated learning, trains successive generations of agents to imitate reconstructions rather than messages, enforcing a tighter communication bottleneck. Third, pairwise distance maximization regularizes message diversity by encouraging high distances between messages, with formal links to entropy maximization. Our method significantly improves both the efficiency and compositionality of the learned messages on the Shapes3D and MPI3D datasets, surpassing prior discrete communication frameworks in both reconstruction accuracy and topographic similarity. This work provides new theoretical and empirical evidence for the emergence of structured, generalizable communication protocols from simplicity-based inductive biases.
@@ 1 Introduction
In natural languages, compositionality enables humans to communicate an infinite number of ideas using a finite set of elements [9]. This principle enables speakers to flexibly combine known words and structures to convey novel meanings, supporting flexible and generalizable communication across diverse and previously unseen contexts.
Works on the emergence of compositionality in language [26, 27, 29] have argued that two opposing bottlenecks are necessary for the natural selection of compositional languages between generations of speakers. On the one hand, successful communication requires high expressivity in order to usefully describe the world. On the other, models of emergent communication such as the iterated language framework [26] state that natural speakers naturally minimize the complexity of languages through cultural transmission, implying that more simple and compressive languages are more easily passed on to new speakers. These opposite pressures of expressivity and compressibility generate a trade-off, highlighted in several works [27, 16, 2], argued to be optimized by compositional languages. When only one of the two pressures is present, languages tend to degrade: excessive compressibility leads to ambiguous, degenerate languages, while prioritizing expressivity alone results in holistic languages without structure, which tend to be more difficult to teach and learn [41].
Inducing compositional behavior in neural networks remains a major challenge [30, 22, 48, 25]. A growing body of work suggests that a model’s ability to generalize compositionally is highly sensitive to its training conditions [31, 10, 49, 32, 11], including factors such as data distribution, learning objectives, and task design. One promising direction is the study of emergent communication, where discrete languages evolve for coordination between independent neural agents.
In this work, we build on the Lewis reconstruction game framework [47] to study emergent communication in a reconstruction task. A sender encodes an image into a discrete message and a receiver reconstructs the original. Drawing on theories of language evolution [26, 27], we develop a novel framework for inducing compositional communication grounded in simplicity bias.
@@
Progressive Decoding, in which the receiver makes reconstruction predictions after each incoming symbol, rather than waiting until the full message is received. This biases the system towards using intermediate reasoning steps, a process shown to support compositional generalization in settings such as chain-of-thought reasoning [56]. While this mechanism does not directly improve reconstruction accuracy, it yields more efficient and structured communication.
Final-state imitation modifies the standard imitation phase in the iterated learning (IL) framework of cultural evolution [27, 28, 29]. Instead of imitating the entire message, the student is trained to reproduce only the final output of the receiver, effectively reducing the information transmitted across generations. This tighter generational bottleneck increases pressure for compressibility in the emergent communication protocol [53], thereby promoting the emergence of compositional structure as a necessary condition for successful transmission. Empirically, this leads to increased compositionality with minimal degradation in reconstruction accuracy.
Pairwise distance maximization as regularization term, i.e. as a further pressure towards increased diversity in the emergent language, that encourages exploration during the imitation phase. This term pushes the student sender to maximize an approximate Hamming distance between messages in a given batch, finding the most "diverse" protocol. We prove that this regularization gives a lower bound for entropy maximization [57], and an upper bound on the contrastive learning loss NT-Xtent [8].
@@ 2.1 Problem setup: recovering compositional representations
In a formalism similar to [44], we model a dataset $\mathcal{D}=\{x=\mathsf{GenX}(\mathbf{G}):\ \mathbf{G}\in\mathcal{G}\}$, i.e. data $x$ are created via a deterministic function $\mathsf{GenX}$ from a set $\mathcal{G}$ of generating factors $\mathbf{G}$. We have access to samples from $\mathcal{D}$ but not to $\mathsf{GenX}$ or to the structure of $\mathbf{G}$, and the goal is to reproduce the set of $x$, i.e., generate images $\widehat{x}$ that approximate the distribution of $x$ in a natural metric such as $\mathbb{E}_{x\in\mathcal{D}}[\text{MSE}(x,\widehat{x})]$, where MSE is the mean squared error on pixels.
@@ 2.2 Approach: Lewis reconstruction game
We frame the above reconstruction problem in the form of a cooperative game between two agents: a sender $S$ and a receiver $R$. The goal of $S$ will be to build the reconstruction factors $\mathbf{G}^{\prime}$ and the receiver $R$ will map the factors to an image $\widehat{x}$. The two agents aim to create a protocol that reconstructs $x$ to good accuracy. This setup fits in the general class of Lewis Reconstruction Games (LRG) [47], a subclass of so-called Lewis Signaling Games (LSGs) [33].
Within the framework of signaling games, the factors $\mathbf{G}^{\prime}$ are interpreted as a language which $S$ uses to communicate $x$ to $R$ as accurately as possible. The allowed message space is finite, i.e. $\mathbf{G^{\prime}}\in\mathcal{M}:=V^{C}$, and we assume that $|V|^{C}\gg|\mathcal{G}|$, so that the size of the message space is not a limiting factor on the emergent communication. The following diagram summarizes our notations:
$\mathcal{G}=\{\mathbf{G}\}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{D}=\mathsf{GenX}(\mathcal{G})\subseteq\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}=V^{C}\stackrel{{\scriptstyle R}}{{\longrightarrow}}\mathcal{X}$
@@ 2.3 Base model architecture
We adopt a standard emergent communication setup as described by Xu et al. [59], see implementation details in Section B.1. Sender and Receiver are implemented using the EGG framework [24]. Input images are encoded into latent representations using a pretrained VAE visual backbone. The sender encodes this latent input into a discrete message using an LSTM and then transmits it through a fixed-length communication channel. Message generation is made differentiable via a Gumbel-Softmax bottleneck [19]. The receiver takes the full message as inputs in an LSTM, and produces a latent vector, which is subsequently decoded into a reconstructed image by the VAE decoder. Reconstruction loss is computed between the original and generated images using Mean Squared Error (MSE). Both VAE encoder and decoder remain frozen during training.
3 Methodology
This section describes our proposed improvements to the EL framework [59] (cf. Figure 1(a)).
@@ 3.1 Progressive Decoding (PD)
Our first proposed enhancement to the baseline communication protocol aims to improve the efficiency of the emergent language determined by Sender and Receiver, by condensing information in fewer message tokens. This type of efficiency has been argued to be necessary for the emergence of natural language properties [21, 36].
Standard reconstruction games typically optimize a reconstruction loss computed only at the end of the communication phase, after the receiver has observed the full message [47], and as shown by Rita et al. [46], without an incentive to efficiently use the information received in the message, the receiver only updates its internal state after the full message has been received. To improve communication efficiency, we propose the following Progressive Decoding (PD) objective: the receiver generates a prediction after each received symbol, and our interaction phase loss explicitly weights the reconstruction error of each sub-message, as follows.
@@
Interpretable sub-messages: The model is encouraged to structure messages such that an interpretation by $R_{\omega}$ is available for sub-messages, making each symbol accountable for partial reconstruction. This echoes findings in prior theoretical and empirical work on chain-of-thought and compositionality [55, 20, 7, 38, 56, 40], which show that forcing intermediate semantic consistency or intermediate reasoning steps encourages models to show increased compositional behavior.
Tuning the efficiency pressure: Small $\lambda$ yield fast but potentially coarse reconstructions, while larger values favor detailed outputs at the cost of longer, possibly redundant messages. In our experiments (see Section 4.4), we explore this trade-off empirically and show that $\lambda>1$ achieves the best performance across generalization, communication efficiency metrics, and compositionality.
@@ 3.2 Imitation Phase
Iterated learning is known to add compressibility pressure to emergent communication schemes [26], forcing languages to be "easy to learn", thus favoring simple communication schemes that can be compressed: this is known as a communication (or generational) bottleneck. It is argued [26] that languages that remain relatively stable even when a learner only observes a small subset of the language of the previous generation need to be compositional.
Our case (and in fact, most deep learning applications of IL) differ from the original IL formulation [26], in that the transmission bottleneck does not constrict the amount of examples presented to the student, but rather constricts the information obtainable from these examples in order to perfectly reconstruct the language. This paradigm shift for transmission bottlenecks is at the heart of discrete bottlenecks [42, 47, 46] and noisy bottlenecks [44].
@@ 3.2.1 Final-state imitation (FiSI)
State space imitation: As opposed to other IL implementations [41, 42, 34], our models are not directed to match the student protocols from different iterations in the message space, but rather in the state space, using the frozen receiver network from the interaction phase. This change allows a wider range of sender strategies, as the new loss is more permissive.
@@
Our loss $\mathcal{L}^{rec}_{\phi^{t+1}}$ only depends on the state associated to the full message, without testing reconstruction on sub-messages. This change allows the intermediate message tokens of student sender $S_{\phi^{t+1}}(x)_{i},i<C$ to drift further away from the ones of the teacher mapping $S^{\star}_{t}(x)_{i}$ and thus more easily find different strategies with the same end result.
State space imitation is justified as follows. Note that as message space is discrete, there exists a value $\delta_{0}$ such that any two messages at distance $<\delta_{0}$ in fact coincide. Then $d_{\mathcal{M}}(S^{\star}_{t}(x),S_{t+1}(x))<\delta_{0}$ always implies $d_{\mathcal{X}}(R^{\star}_{t}(S^{\star}_{t}(x)),R^{\star}_{t}(S_{t+1}(x)))=0$, while the vice-versa is not always true: if $R^{\star}_{t}$ is not injective, any $S_{t+1}(x)\in(R^{\star}_{t})^{-1}(S^{\star}_{t}(x))$ will preserve zero distance of $R^{\star}_{t}$-images. If $R^{\star}_{t}$ has mild Lipschitz regularity assumption, this principle extends to the regime of small loss, showing that our new loss is less restrictive on the choices of $S_{t+1}$ than the traditional message space loss. Thus, state space imitation enables wider sender strategy exploration without penalization losses.
@@
Figure 1: (a) Overview of our proposed architecture. Interaction Phase: The sender $S_{\phi^{t}}$ and receiver $R_{\omega^{t}}$ are jointly trained to minimize the reconstruction error between the input state $x$ and the predicted states $\{x^{\prime}\}$, by encoding $x$ into a message $m^{t}_{x}$. Imitation Phase: A new sender $S_{\phi^{t+1}}$ is trained to imitate the final predicted output of $R_{\omega^{t}}(m^{t}_{x})$, while also maximizing pairwise message diversity to encourage exploration. (b) Qualitative example of image decoding: the receiver reconstructs the input image from the sender’s message at the final decoding step.
The choice of using just final states for reconstruction is justified using the following result, which states that final state reconstruction allows greater freedom during imitation for the student sender.
Proposition 3.1 (See Prop. E.1).
Set $d_{i}^{S}(x):=d_{\mathcal{X}}(R^{\star}(S^{\star}(x)_{[i]}),R^{\star}(S(x)_{[i]}))$ and for $\epsilon>0$ set
1. For $i<C$ and any choice of $M,\epsilon>0$ we have $\mathcal{R}_{[i]}(\mathcal{S}_{full}^{M})\subset\mathcal{R}_{[i]}(\mathcal{S}_{final}^{\epsilon})$ where $\mathcal{R}_{[i]}$ is the restriction of values of $S$ to sub-messages of length $i$, $\mathcal{R}_{[i]}S(x):=S(x)_{[i]}$.
2. In general, $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{\epsilon/C}$. If $\mathbb{E}_{x}[d_{j}^{S}(x)]$ is non-decreasing in $j$ then $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{\epsilon}$ as well.
The freedom of exploration ensured by item 1 of the proposition is our main motivation. By only constraining the reconstruction of the full message, we allow much more freedom for the encoding strategy of $S(x)$ at earlier sub-messages at each iteration.
Item 2 of Prop. 3.1 shows that the two losses in the definition of $\mathcal{S}_{full},\mathcal{S}_{final}$ give comparable guarantees. At the beginning of iterations we will need to use the general bound with $\epsilon/C$, and as $R^{\star}$ will have been trained on more and more messages during the iteration, the hypothesis of $\mathbb{E}_{x}[d_{i}^{S}(x)]$ non-increasing will become true allowing for the sharper bound with the same value of $\epsilon$, as $R^{\star}$ (an LSTM taking as input successive message tokens) will become efficient in using sub-messages of pairs $(m,m^{\prime})=(S(x),S^{\star}(x))$ to gradually distill information allowing to distinguish them for a large set of $S,x$.
3.2.2 Pairwise distance maximization (PDM) between messages
To further improve protocol diversity and prevent message collapse during the imitation phase, we introduce a regularization objective that promotes dissimilarity between messages within each batch. In addition to enabling a larger search space for optimal student strategies. Enhancing the diversity of student strategies also fits the general IL argument according to which compositional languages are the most resistant to noisy language transmission [26, 29, 28].

### v4

@@ Abstract
Compositionality in knowledge and language—the ability to represent complex concepts as a combination of simpler ones—is a hallmark of human cognition and communication. Despite recent advances, deep neural networks still struggle to acquire this property reliably. Neural models for emergent communication look to endow artificial agents with compositional language by simulating the pressures that form human language. In this work, we introduce CELEBI [^1] Our oficial implementation can be found at https://github.com/SugarFreeManatee/CELEBI (Compressive-Expressive Language Emergence through a discrete Bottleneck and Iterated learning), a novel self-supervised framework for inducing compositional representations through a reconstruction-based communication game between a sender and a receiver. Building on theories of language emergence and the iterated learning framework, we integrate three mechanisms that jointly promote compressibility, expressivity, and efficiency in the emergent language. First, Progressive Decoding incentivizes intermediate reasoning by requiring the receiver to produce partial reconstructions after each symbol. Second, Final-State Imitation trains successive generations of agents to imitate reconstructions rather than messages, enforcing a tighter communication bottleneck. Third, Pairwise Distance Maximization regularizes message diversity by encouraging high distances between messages, with formal links to entropy maximization. Our method significantly improves both the efficiency and compositionality of the learned messages on the Shapes3D and MPI3D datasets, surpassing prior discrete communication frameworks in both reconstruction accuracy and topographic similarity. This work provides new theoretical and empirical evidence for the emergence of structured, generalizable communication protocols from simplicity-based inductive biases.
@@ 1 Introduction
In natural languages, compositionality enables humans to communicate an infinite number of ideas using a finite set of elements [chomsky2014aspects-of-syntax]. This principle enables speakers to flexibly combine known words and structures to convey novel meanings, supporting flexible and generalizable communication across diverse and previously unseen contexts.
Works on the emergence of compositionality in language [iteratedlearningkirby, Kirby2002, KIRBY201587] have argued that opposing pressures are necessary for the natural selection of compositional languages between generations of speakers [elmoznino2024complexitybasedtheorycompositionality, beppu2009iterated]. On the one hand, successful communication requires high expressivity in order to usefully describe the world, allowing speakers to produce distinct messages for a wide range of meanings. On the other hand, models of emergent communication such as the iterated learning framework [iteratedlearningkirby] state that natural speakers tend to minimize the complexity of languages through cultural transmission, implying that simpler and more compressive languages are more easily passed on to new speakers.
These opposite pressures of expressivity and compressibility thus generate a trade-off, highlighted in several works [Kirby2002, elmoznino2024complexitybasedtheorycompositionality, beppu2009iterated] which is argued to be optimized by compositional languages, whereby languages evolve to maximize communicative efficiency – remaining expressive enough to convey diverse meanings while being simple and structured enough to be easily learned and transmitted by successive generations of speakers. When either pressure is dominant, undesirable languages tend to emerge: (a) excessive language compression leads to degenerate languages where multiple meanings are mapped to the same messages making them ambiguous and thus hard to use, and (b) prioritizing expressivity alone produces in holistic languages i.e. languages where messages are mapped arbitrarily to meanings without respecting their structures, hindering their transmission across generations [ren2020Compositional].
Enabling machines to generalize compositionally is thought to be crucial for them to quickly adapt to novel situations beyond their training experience [lake2017building]. Inducing compositional behavior in neural networks remains a major challenge [lake2017original-scan, keysers2019cfq, ruis2020gscan, kim2020cogs]. A growing body of work suggests that a model’s ability to generalize compositionally is highly sensitive to its training conditions [lake2019meta-seq2seq, conklin2021meta, ruis2022improving, lake2023human, delrio2025data], including factors such as data distribution, learning objectives, and task design. One promising direction is the study of emergent communication, where discrete languages evolve for coordination between independent neural agents.
In this work, we build on the Lewis reconstruction game framework [rita2022emergent] to study emergent communication in a reconstruction task. A sender encodes an image into a discrete message and a receiver reconstructs the original. Drawing on theories of language evolution [iteratedlearningkirby, Kirby2002], we develop a novel framework for inducing compositional communication grounded in simplicity bias.
@@
Progressive Decoding, in which the receiver makes reconstruction predictions after each incoming symbol, rather than waiting until the full message is received. This biases the system towards using intermediate reasoning steps, thereby imposing a pressure towards lower-complexity and less holistic encodings. While this mechanism does not directly improve reconstruction accuracy, it yields more efficient and structured communication.
Final-State Imitation modifies the standard imitation phase in the iterated learning (IL) framework of cultural evolution [Kirby2002, kirby2014iterated, KIRBY201587]. Instead of imitating the entire message, the student is trained to reproduce only the final output of the receiver, effectively reducing the information transmitted across generations. This tighter generational bottleneck increases pressure for compressibility in the emergent communication protocol [smithiterated], thereby promoting the emergence of compositional structure as a necessary condition for successful transmission. Empirically, this leads to increased compositionality with minimal degradation in reconstruction accuracy.
Pairwise Distance Maximization as a regularization term, i.e., as a further pressure towards increased diversity in the emergent language, that encourages exploration during the imitation phase. This term pushes the student sender to maximize an approximate Hamming distance between messages in a given batch, finding the most "diverse" protocol. We prove that this regularization gives a lower bound for entropy maximization [entropymax], and an upper bound on the contrastive learning loss NT-Xtent [chen2020simpleframeworkcontrastivelearning].
The opposing pressures induced by these three mechanisms impose a tight regularization on the complexity of the emergent language. These ingredients make explicit the pressures proposed in the cognitive science and linguistic literature to shape language evolution, thus aligning with the IL paradigm and leading to measurably higher compositionality.
@@ 2.1 Problem setup: recovering compositional representations
[…]
@@ 2.2 Approach: Lewis reconstruction game
We frame the above reconstruction problem in the form of a cooperative game between two agents: a sender $S$ and a receiver $R$. The goal of $S$ will be to build the reconstruction factors $\mathbf{G}^{\prime}$ and the receiver $R$ will map the factors to an image $\widehat{x}$. The two agents aim to create a protocol that reconstructs $x$ to good accuracy. This setup fits in the general class of Lewis Reconstruction Games (LRG) [rita2022emergent], a subclass of so-called Lewis Signaling Games (LSGs) [lewis1969convention].
Within the framework of signaling games, the factors $\mathbf{G}^{\prime}$ are interpreted as a language which $S$ uses to communicate $x$ to $R$ as accurately as possible. The allowed message space is finite, i.e. $\mathbf{G^{\prime}}\in\mathcal{M}:=V^{C}$, and we take $C$ fixed and large enough that $|V|^{C}\gg|\mathcal{G}|$, so that the size of the message space is not a limiting factor on the emergent communication. The following diagram summarizes our notations:
$\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{D}\subseteq\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}=V^{C}\stackrel{{\scriptstyle R}}{{\longrightarrow}}\mathcal{X}$
@@ 2.3 Base model architecture
We adopt a standard emergent communication setup as described by elvae, see implementation details in § B.1. Sender and Receiver are implemented using the EGG framework [kharitonov:etal:2021]. Input images are encoded into latent representations using a pretrained VAE visual backbone. The sender encodes this latent input into a discrete message using an LSTM and then transmits it through a fixed-length communication channel. Message generation is made differentiable via a Gumbel-Softmax bottleneck [jang2017categoricalreparameterizationgumbelsoftmax]. The receiver takes the full message as inputs in an LSTM, and produces a latent vector, which is subsequently decoded into a reconstructed image by the VAE decoder. Reconstruction loss is computed between the original and generated images using Mean Squared Error (MSE). Both VAE encoder and decoder remain frozen during training.
3 Methods
This section describes our proposed improvements to the EL framework [elvae] (cf. Figure 1(a)). Intuitively, our additions attempt to increase the language drift between iterations of speakers, in order to thoroughly explore the landscape of possible languages, while maintaining efficient and useful communication. For the latter, we introduce Progressive Decoding (PD), to serve as an anchor for the emergent language during interaction, and for the former we introduce Final-State Imitation (FiSI) and Parwise Distance Maximization (PDM), which aim to incentivize exploration during the imitation phase. We expect the combination of these factors to show the strongest impact towards compositionality in the learned language, more so than the factors taken separately.
@@ 3.1 Progressive Decoding (PD)
Our first proposed enhancement to the baseline communication protocol aims to improve the efficiency of the emergent language determined by the Sender and Receiver, by condensing information in fewer message tokens. This type of efficiency has been argued to be necessary for the emergence of natural language properties [Kanwal2017-hs, Luna2020InternalAE].
Standard reconstruction games typically optimize a reconstruction loss computed only at the end of the communication phase, after the receiver has observed the full message [rita2022emergent], and as shown by rita-etal-2020-lazimpa, without an incentive to efficiently use the information received in the message, the receiver only updates its internal state after the full message has been received. To improve communication efficiency, we propose the following Progressive Decoding (PD) objective: the receiver generates a prediction after each received symbol, and our interaction phase loss explicitly weights the reconstruction error of each sub-message, as follows.
@@
Interpretable sub-messages: The model is encouraged to structure messages such that an interpretation by $R_{\omega}$ is available for sub-messages, making each symbol accountable for partial reconstruction. This echoes findings in prior theoretical and empirical work on chain-of-thought and compositionality [velickovic2020Neural, jiang2021inducing, chakravarthy2022systematicity, nye2022show, wei2022chain, prystawski2024think], which show that forcing intermediate semantic consistency or intermediate reasoning steps encourages models to show increased compositional behavior.
Tuning the efficiency pressure: Small $\lambda$ yields fast but potentially coarse reconstructions, while larger values favor detailed outputs at the cost of longer, possibly redundant messages, and in the limit $\lambda\to\infty$ (1) becomes equivalent to using just the last term in brackets, i.e., the full-length message reconstruction. In our experiments (see Section 5.2), we explore this trade-off empirically and show that $\lambda\equiv 1.5$ achieves the best performance across generalization, communication efficiency metrics, and compositionality.
@@ 3.2 Imitation Phase
Iterated learning is known to add compressibility pressure to emergent communication schemes [iteratedlearningkirby], resulting in languages that are "easy to learn", thus favoring simple communication schemes that can be compressed: this is known as a communication (or generational) bottleneck. It is argued [iteratedlearningkirby] that languages that remain relatively stable even when a learner only observes a small subset of the language of the previous generation need to be compositional.
Our case (and in fact, most deep learning applications of IL) differ from the original IL formulation [iteratedlearningkirby], in that the transmission bottleneck does not constrict the amount of examples presented to the student, but rather constricts the information obtainable from these examples in order to perfectly reconstruct the language. This paradigm shift for transmission bottlenecks is at the heart of discrete bottlenecks [ren2020Compositional, rita2022emergent, rita-etal-2020-lazimpa] and noisy bottlenecks [ren2024improving].
@@ 3.2.1 Final-State Imitation (FiSI)
State space imitation: As opposed to other IL implementations [ren2020Compositional, ren2020Compositional, li2019easeofteachinglanguagestructureemergent], our models are not directed to match the student protocols from different iterations in the message space, but rather in the state space, using the frozen receiver network from the interaction phase. This change allows a wider range of sender strategies, as the new loss is more permissive.
@@
Our loss $\mathcal{L}^{rec}_{\phi^{t+1}}$ only depends on the state associated to the full message, without testing reconstruction on sub-messages. This change allows the intermediate message tokens of the student sender $S_{\phi^{t+1}}(x)_{i},i<C$ to drift further away from the ones of the teacher mapping $S^{\star}_{t}(x)_{i}$ and thus more easily find different strategies with the same end result.
State space imitation is justified as follows. Note that as message space is discrete, there exists a value $\delta_{0}$ such that any two messages at distance $\delta_{0}$ in fact coincide. Then $d_{\mathcal{M}}(S^{\star}_{t}(x),S_{t+1}(x))<\delta_{0}$ always implies $d_{\mathcal{X}}(R^{\star}_{t}(S^{\star}_{t}(x)),R^{\star}_{t}(S_{t+1}(x)))=0$, while the vice-versa is not always true: if $R^{\star}_{t}$ is not injective, any $S_{t+1}(x)\in(R^{\star}_{t})^{-1}(S^{\star}_{t}(x))$ will preserve zero distance of $R^{\star}_{t}$-images. If $R^{\star}_{t}$ has mild Lipschitz regularity assumption, this principle extends to the regime of small loss, showing that our new loss is less restrictive on the choices of $S_{t+1}$ than the traditional message space loss. Thus, state space imitation enables wider sender strategy exploration without penalization losses.
@@
Figure 1: (a) Overview of our proposed architecture. Interaction Phase: The sender $S_{\phi^{t}}$ and receiver $R_{\omega^{t}}$ are jointly trained to minimize the reconstruction error between the input state $x$ and the predicted states $\{x^{\prime}\}$, by encoding $x$ into a message $m^{t}_{x}$. Imitation Phase: A new sender $S_{\phi^{t+1}}$ is trained to imitate the final predicted output of $R_{\omega^{t}}(m^{t}_{x})$, while also maximizing pairwise message diversity to encourage exploration. (b) Qualitative example of image decoding: the receiver reconstructs the input image from the sender’s message for each sub-message, giving a series of reconstructions. The useful length of a message corresponds to the first reconstruction that has error below a threshold $\epsilon>0$: in this example the last two message tokens are not useful in that they do not add to the reconstruction accuracy.
The choice of using just final states is justified because final state reconstruction allows greater freedom during imitation for the student sender. By only constraining the reconstruction of the full message, we allow much more freedom for the encoding strategy of $S(x)$ at earlier sub-messages at each iteration (see Proposition E.1 for a formal statement and proof).
3.2.2 Pairwise Distance Maximization (PDM)
To further improve protocol diversity and promote exploration of the language space during the imitation phase, we introduce a regularization objective that promotes dissimilarity between messages within each batch. In addition to enabling a larger search space for optimal student strategies. Enhancing the diversity of student strategies also fits the general IL argument according to which compositional languages are the most resistant to noisy language transmission [iteratedlearningkirby, KIRBY201587, kirby2014iterated].